### v1

# Class Distance Weighted Cross Entropy Loss for Classification of Disease Severity

Abstract
Assessing disease severity involving ordinal classes, where each class represents increasing levels of severity, benefit from loss functions that account for this ordinal structure. Traditional categorical loss functions, like Cross-Entropy (CE), often perform suboptimally in these scenarios. To address this, we propose a novel loss function, Class Distance Weighted Cross-Entropy (CDW-CE), which penalizes misclassifications more harshly when classes are farther apart. We evaluated CDW-CE on the Labeled Images for Ulcerative Colitis (LIMUC) dataset using various deep architectures. Its performance was compared against several categorical and ordinal loss functions. To analyze the quality of latent representations, we used t-distributed stochastic neighbor embedding (t-SNE) visualizations and quantified their clustering with the Silhouette Score. We also compared Class Activation Maps (CAM) generated by models trained with CDW-CE and CE loss, incorporating domain expert feedback to evaluate alignment with expert knowledge. Our results show that CDW-CE consistently improves performance in ordinal image classification tasks. It achieves higher Silhouette Scores, indicating better differentiation of class representations, and its CAM visualizations demonstrate a stronger focus on clinically significant regions, as confirmed by domain experts.
keywords:
ordinal classification; convolutional neural network; computer-aided diagnosis; loss function
Introduction
In ordinal classification tasks, it is common to use the distance relation between different classes. This relation is evident in many examples, including disease severity classification tasks where classes can correspond to the severity of disease, and their ordered nature can be exploited with custom loss functions [1, 2, 3, 4], instead of employing categorical loss functions.
In the medical domain, disease classes are assigned according to the symptom severity, thus implying an ordinal relation between classes[5]. Misclassification between further separated classes is more detrimental than the misclassification of relatively closer classes. For example, in the context of a medical diagnosis tool, erroneously categorizing a patient as healthy when they have a severe condition is considerably more problematic than misclassification between a mild or moderate condition. On the other hand, commonly used loss functions consider each class in isolation and do not account for any distance information between them, and both misclassification cases will have the same impact on the error. Consequently, both types of misclassifications are accorded equal weight in terms of error, disregarding the practical impact. To address this, a possible approach is mathematically representing the distance information in a loss function, thereby penalizing misclassifications between classes that are further apart more severely than those closer.
In our earlier work, we investigated application of ordinal loss functions for ulcerative colitis severity estimation [6]. In this work, we extend the proposed loss function, Class Distance Weighted Cross Entropy (CDW-CE) and demonstrate its capabilities. In addition, we expand the CDW-CE by introducing an additive margin term into the loss function, analyzing class discrimination properties of models trained with different loss functions, and demonstrating its effectiveness with supplementary experiments comparing against various loss functions in the literature.
Ulcerative Colitis is a disease for which several deep learning-based models have been proposed to classify disease severity. In this problem, disease severity is represented in increasing sequential order, and classification models are trained to predict the disease severity classes corresponding to input images correctly. Labeled Images for Ulcerative Colitis (LIMUC) [7] dataset is a publicly available dataset with ordinal classes and imbalanced class distributions. We have used this dataset to comparatively evaluate CDW-CE and other standard loss functions namely Cross-Entropy (CE), Mean Squared Error (MSE), CORN framework [2], cross-entropy with an ordinal loss term (CO2) [4], and ordinal entropy loss (HO2) [4] from the literature.
We leveraged t-distributed stochastic neighbor embedding (t-SNE) plots to analyze the latent representations of models trained with CDW-CE and CE. We have also calculated the Silhouette Score of the t-SNE plots to assess the latent representation quality of different methods. Class Activation Map (CAM) outputs from CDW-CE and CE methods and raw endoscopic images are presented to medical experts, eliciting their opinions on the alignment of CAM outputs with observed symptoms.
The main contribution of this paper is the proposal of an advanced ordinal loss function, Class Distance Weighted Cross-Entropy (CDW-CE), with hyperparameters of $\alpha$, the class distance penalization power term, and margin, the separation robustness enforcer. As detailed in the experimental evaluation section, these hyperparameters can be tuned to achieve better results.
Related Work
Ordinal categories are frequently observed in real-world situations [8], particularly in the medical domain [9], such as disease severity classifications [10]. For ordinal classification problems, tailored methodologies, including novel loss functions that take ordinality into account, are reported to outperform traditional methods [11]. This section provides an overview of methods making use of such loss functions.
An end-to-end deep learning method for ordinal regression, focusing on age estimation from facial images, was introduced in [12]. This approach involved transforming ordinal regression into a series of binary classification sub-problems and employing a multiple-output CNN for joint task resolution. By using $N-1$ output nodes for $N$ classes and applying label extension to the target class, their method is reported to perform better than other ordinal regression techniques, such as metric learning. However, this approach exhibited inconsistencies in output ranking within the subtasks. Consistent Rank Logits (CORAL) framework [1] addresses rank-inconsistencies in extended binary classification models. This framework tackles the problem of inconsistent classifier predictions in neural networks, where model confidence should ideally decrease as the rank increases. CORAL differs from the method proposed in [12], employing weight sharing in the penultimate layer during training, ensuring classifier consistency. The authors demonstrated the effectiveness of CORAL when integrated into common CNN architectures like ResNet, resulting in enhanced predictive performance for age estimation tasks. Conditional Ordinal Regression for Neural Network (CORN) framework [2] relaxes the constraints on the penultimate layer of the CORAL framework and incorporates conditional probabilities. Experiments on various datasets, including MORPH-2, AFAD, AES, and FIREMAN, showed that CORN outperformed previous methods. However, as mentioned earlier, these techniques need modification of both model architecture and labeling structure. An alternative approach for ordinal classification is integrating uni-modality distribution into the model’s output predictions, penalizing inconsistencies in posterior probability distribution between neighboring labels alongside the primary loss function, often using cross-entropy.
Non-parametric ordinal loss [3] encourages output probabilities to follow an unimodal distribution by penalizing inconsistent probability pairs in the output. This approach was validated on various tasks, including breast cancer grading, predicting the decade of a photograph, and age estimation. An extension of this technique by adding unimodality losses to cross-entropy (CO2) and entropy losses (HO2) for the final loss function was proposed in [4], and it was validated using the Herlev dataset [13], consisting of cervical cell images, and various CNN architectures. Both studies reported improved performance compared to the method in [12] and cross-entropy loss.
CO2 and HO2 losses are frequently used for comparison in this study. In these equations, CE and H refer to Cross entropy and Entropy losses, respectively, $y_{n}$ is ground truth, $\hat{y}_{n}$ is model predictions, $K$ is the total number of classes, $\lambda$ is a parameter that determines the strength of the unimodal loss, and $\delta$ is the margin term.
An alternative approach involves using regression to estimate a single continuous value at the output or applying a sigmoid activation function to constrain predictions within the range of [0, 1], which are then transformed into discrete levels using thresholds or probability distributions. Beckham et al. [14] proposed adding a single-node layer on top of the final layer with squared-error loss, suggesting that processing the final output through a sigmoid function followed by multiplication by (K - 1) yields better results, with rounding to the nearest integer for inference. They reported improved performance compared to standard cross-entropy loss. Polat et al. [15] developed a regression-based deep learning system, leveraging convolutional neural networks for enhanced accuracy and robustness in classifying Ulcerative Colitis severity. However, It is worth noting that regression-based approaches have generally been reported to exhibit inferior performance in various studies [4], [3].
Methods
This work introduces an ordinal loss function designed to improve model performance in ordinal classification tasks, as previously discussed. This section provides detailed information on developing our custom loss function, the experimental setup, and the dataset used.
The main idea behind employing an ordinal loss function is demonstrated in Figure 1. Unlike categorical loss functions, such as cross-entropy loss, which treat bimodal and unimodal distributions the same, an ordinal loss function can address issues specific to ordinal datasets. Mayo Endoscopic Scoring (MES) is used to classify the severity of Ulcerative Colitis, an inflammatory bowel disease. MES classes are ordinal, and thus, the misclassification of MES-2 (moderate disease severity) as MES-0 (no disease, healthy) is comparatively more harmful than the misclassification of adjacent classes (moderate and mild disease severities). We aim to encourage deep learning models to predict classes correctly while resembling an unimodal distribution to improve the robustness of the predictions.
Figure 1: Assuming MES-2 is the true class, the loss calculated with CE is the same for both cases. On the other hand, the unimodal distribution is more intuitive for ordinal classification.
Class Distance Weighted Cross Entropy Loss
We start with the standard cross-entropy loss (Eq. 1) as the baseline of our method.
$\textrm{CE}=-\sum_{i=0}^{N-1}{y_{i}\times\log{\hat{y}_{i}}}=-\log{\hat{y}_{c}}$ (1)
In Equation 1, $i$ refers to the index of the class in the output layer, $c$ is the index of the ground-truth class, $y$ is the ground-truth label, and $\hat{y}$ refers to the prediction. Since one-hot encoding is used for ground-truth labels, $y_{i}$ becomes 00 for $\forall i\neq c$; therefore, CE loss only calculates the negative log-likelihood of the confidence of the ground-truth class (This is why CE is sometimes called the negative log-likelihood function).
The proposed Class Distance Weighted Cross Entropy (CDW-CE) (Eq. 2) is a non-parametric loss function for multi-class classification. This function penalizes misclassifications more severely when they diverge farther from the true class. The addition of a class distance penalty term aims to regulate the penalty of class misprediction based on the distance between the classes.
Moreover, including a margin term encourages tighter grouping within the same class while enhancing separation between different classes. CDW-CE comprises two main parameters: a logarithmic loss term indicating the extent of deviation for each misclassification from the correct value and a coefficient term that considers the distance to the ground-truth class. The parameter $\alpha$ in the loss coefficient governs the intensity of the penalty, where higher values lead to harsher penalties for misclassifications that significantly deviate from the true class.
[…]
CDW-CE loss is differentiable and can be used directly in backpropagation calculations. As shown in Equation 3, the derivative term increases when the prediction value deviates further from its target value of zero.
$\frac{d(L)}{d(\hat{y}_{i})}=\frac{|i-c|^{\alpha}}{1-\hat{y}_{i}}$ (3)
Adding a margin to the loss function enforces intra-class compactness and increases inter-class distances, as shown in ArcFace [16], CosFace [17], SphereFace [18], and NormFace [19]. This is incorporated into CDW-CE loss by adding a margin to the class probabilities. CDW-CE with margin loss (Equation 4) is expected to improve the overall classification performance. However, this performance improvement comes at the cost of increased hyperparameter tuning and training time, as the margin term needs to be tuned for the specific dataset and task.
[…]
Experimental Evaluation
We used LIMUC [7] dataset and employed ResNet18 [20], Inception-v3 [21], and MobileNet-v3-large [22] architectures for experimental evaluation. In addition to the classification performance metrics, we have also calculated remission scores, a score used by medical experts that bins lower MES classes of 0 and 1 and higher MES classes of 2 and 3 together. This remission score is essential for domain experts aiming to classify whether a case is healthy or contains mild symptoms, which corresponds to the remission class, or contains moderate to severe symptoms, which is the non-remission class. We report the comparative performance evaluation of various methods using Quadratic Weighted Kappa (QWK) for multi-class classification, Kappa for binary (remission) classification, and accuracy, precision, recall, and F1 scores.
The proposed ordinal loss function Class Distance Weighted Cross Entropy (CDW-CE) evaluates the performance of a given prediction according to the ordinal nature and penalizes predictions further from the ground truth. The degree of penalization is adjustable through the $\alpha$ hyperparameter, which is a power term above the class distance coefficient. We analyzed the impact of $\alpha$ by varying it across integer values within the range of 1 to 10. For each model (ResNet18, Inception-v3, and MobileNet-v3-large), we found the best-performing power term values as 5, 6, and 7, respectively. These experiments are summarized in the Figure 2.
Figure 2: Effects of $\alpha$ for three different models on the mean and standard deviation of QWK scores.
We have implemented CDW-CE and other loss functions from the literature such as Cross-Entropy (CE), Mean Squared Error (MSE), CORN framework [2], cross-entropy with an ordinal loss term (CO2) [4], and ordinal entropy loss (HO2) [4]. We used Cross Entropy (CE) as our baseline and experimentally evaluated the CDW-CE against other loss functions.
The explainability of deep learning models is crucial because they are often deemed black-box systems, given that their features are autonomously learned through an iterative training process. CAMs can be used to visualize and explain where the attention of deep learning models is on the images to overcome this problem. Given a sample input image, CAM demonstrates a deep network’s neural activation map in 2-dimensional. We have obtained CAMs of CDW-CE and CE methods and raw endoscopic images and shared them with the medical experts to get their opinion on which model’s CAM output is more aligned with symptoms. Figure 3 provides a sample of CAM visuals.
Figure 3: A sample CAM output of two different models, trained with and without CDW-CE, presented to medical experts for their feedback.
Dataset
We conducted our experiments on the LIMUC [7] dataset, which comprises endoscopic images with varying severity of Ulcerative Colitis (UC), ranked in increasing order from zero to three. UC is one of the two types of inflammatory bowel disease (IBD). Domain experts, i.e., medical doctors, classify whether the disease is present and rate its severity. The corresponding classes of healthy, mild, moderate, and severe disease conditions align with the ordinal levels of Mayo Endoscopic Scoring (MES) of zero to three. While there are several scoring systems for assessing UC from endoscopic images, the Mayo Endoscopic Score (MES) is the most widely used one [23]. MES assigns an ordinal label ranging from 0 (healthy instance) to 3 (severe symptoms) for symptom severity assessment of UC. Medical experts also employ a remission state scoring system, consisting of two classes, to distinguish between the presence (MES 2 and 3) and absence (MES 0 and 1) of UC.
LIMUC dataset is the largest publicly available UC dataset, comprising 11276 images from 1043 colonoscopy procedures on 564 patients. MES classes from 0 to 3 have 6105, 3052, 1254, and 865 images, respectively. The imbalance of the dataset is evident from the percentages corresponding to 54.1%, 27.1%, 11.1%, and 7.7%, respectively, for each class.
Results
We performed experiments for Ulcerative Colitis classification using the dataset detailed in the previous section. The performance results are complemented by an explainability analysis utilizing t-SNE plots. Additionally, we gathered feedback from domain experts to evaluate whether our loss function contributes to the model’s Class Activation Map (CAM) in better aligning with professional opinion.
Assessment of Ulcerative Colitis Classification Performance
We have employed three deep learning models and six different loss functions, reporting results against various loss functions. The performance results are reported in Table 1 and 2 for multi-class (MES) classification and binary score (remission) classification tasks, respectively.
Table 1: Experiment results for all MES.
Loss Function ResNet18 Inception-v3 MobileNet-v3-Large
QWK Cross-Entropy 0.8296 ± 0.014 0.8360 ± 0.011 0.8302 ± 0.015
MSE 0.8540 ± 0.007 0.8517 ± 0.007 0.8467 ± 0.005
CORN 0.8366 ± 0.007 0.8431 ± 0.009 0.8412 ± 0.010
CO2 0.8394 ± 0.009 0.8482 ± 0.009 0.8354 ± 0.009
HO2 0.8446 ± 0.007 0.8458 ± 0.010 0.8378 ± 0.007
CDW-CE 0.8568 ± 0.010 0.8678 ± 0.006 0.8588 ± 0.006
F1 Cross-Entropy 0.6720 ± 0.026 0.6829 ± 0.023 0.6668 ± 0.028
MSE 0.6925 ± 0.015 0.6881 ± 0.013 0.6946 ± 0.011
CORN 0.6809 ± 0.014 0.6832 ± 0.013 0.6847 ± 0.020
CO2 0.6782 ± 0.014 0.6846 ± 0.016 0.6793 ± 0.012
HO2 0.6856 ± 0.016 0.6901 ± 0.008 0.6741 ± 0.030
CDW-CE 0.7055 ± 0.021 0.7261 ± 0.015 0.7254 ± 0.010
Accuracy Cross-Entropy 0.7566 ± 0.015 0.7600 ± 0.012 0.7564 ± 0.011
MSE 0.7702 ± 0.009 0.7690 ± 0.008 0.7677 ± 0.009
CORN 0.7591 ± 0.009 0.7600 ± 0.008 0.7613 ± 0.012
CO2 0.7601 ± 0.008 0.7654 ± 0.008 0.7572 ± 0.009
HO2 0.7625 ± 0.011 0.766 ± 0.010 0.7583 ± 0.005
CDW-CE 0.7740 ± 0.011 0.7880 ± 0.011 0.7759 ± 0.010
MAE Cross-Entropy 0.2581 ± 0.018 0.2526 ± 0.013 0.2563 ± 0.012
MSE 0.2346 ± 0.009 0.2359 ± 0.009 0.2383 ± 0.009
CORN 0.2517 ± 0.012 0.2497 ± 0.010 0.2480 ± 0.012
CO2 0.2497 ± 0.011 0.2404 ± 0.008 0.2524 ± 0.010
HO2 0.2460 ± 0.011 0.2424 ± 0.011 0.2487 ± 0.005
CDW-CE 0.2300 ± 0.011 0.2147 ± 0.010 0.2272 ± 0.011
The CDW-CE loss function outperforms all other functions for MES classification by achieving top scores in QWK, F1-Score, accuracy, and MAE metrics. On the other hand, the baseline CE loss function is the least performing among all models. MSE, CORN, CO2, and HO2 consistently outperform CE but are inferior to CDW-CE across various CNN models, with varying performance rankings among themselves.
Classifying disease severity into remission and non-remission holds clinical significance. Consequently, we mapped the results in Table 2 into remission classification without training a new model. CDW-CE loss function demonstrates the best performance by achieving the highest Kappa, F1-Score, and accuracy metrics. Contrary to multi-class classification, the CE loss function often performs worst among other loss functions while achieving better scores than CORN and CO2 methods for binary classification. Similar to multi-class classification, MSE exhibits better performance than other loss functions (CORN, CO2, HO2, and CE).
Table 2: Experiment results for remission classification.
Loss Function ResNet18 Inception-v3 MobileNet-v3-Large
Kappa Cross-Entropy 0.8077 ± 0.023 0.8074 ± 0.021 0.8122 ± 0.018
MSE 0.8406 ± 0.013 0.8404 ± 0.017 0.8339 ± 0.012
CORN 0.8191 ± 0.021 0.8077 ± 0.022 0.8203 ± 0.016
CO2 0.8185 ± 0.020 0.8243 ± 0.011 0.8067 ± 0.020
HO2 0.8318 ± 0.015 0.8251 ± 0.015 0.8283 ± 0.018
CDW-CE 0.8521 ± 0.016 0.8598 ± 0.012 0.8592 ± 0.012
F1 Cross-Entropy 0.8419 ± 0.018 0.8420 ± 0.017 0.8451 ± 0.016
MSE 0.8691 ± 0.011 0.8686 ± 0.014 0.8634 ± 0.010
CORN 0.8511 ± 0.016 0.8425 ± 0.018 0.8523 ± 0.013
CO2 0.8513 ± 0.015 0.8561 ± 0.009 0.8404 ± 0.017
HO2 0.8618 ± 0.012 0.8565 ± 0.011 0.8583 ± 0.015
CDW-CE 0.8785 ± 0.013 0.8847 ± 0.010 0.8842 ± 0.010
Accuracy Cross-Entropy 0.9436 ± 0.009 0.9432 ± 0.007 0.9456 ± 0.005
MSE 0.9531 ± 0.004 0.9536 ± 0.006 0.9514 ± 0.005
CORN 0.9473 ± 0.007 0.9429 ± 0.008 0.9473 ± 0.006
CO2 0.9461 ± 0.008 0.9479 ± 0.004 0.9444 ± 0.006
HO2 0.9507 ± 0.005 0.9485 ± 0.005 0.9504 ± 0.005
CDW-CE 0.9566 ± 0.005 0.9590 ± 0.003 0.9588 ± 0.005
The confusion matrices for multi-class and binary classification tasks are shown in Figures 4(a) and 4(b), respectively. Figure 4(a) demonstrates that the CDW-CE loss function decreases mispredictions with a distance larger than one. With decreased mispredictions for two and three-class distances, the loss function manipulates the model learning as expected. Since CDW-CE clusters mispredictions around 1-class distance, naturally, this is also reflected in the better classification as shown in Figure 4(b).
(a) Mean confusion matrix for full MES classification.
(b) Mean confusion matrix for remission classification.
Figure 4: Mean confusion matrix of each CNN model trained with CE and CDW-CE for full MES classification.
Discussion
Explainability Analysis using t-SNE plots
In this section, we use t-SNE plots [24] to visually assess the feature extraction capability of our method by examining high-dimensional data. t-SNE effectively reveals clusters and latent patterns in the data. To complement this, we analyze silhouette scores, which quantify clustering quality by measuring intra-class cohesion and inter-class separation. A higher silhouette score indicates better clustering. While t-SNE plots are helpful for visual insights, silhouette scores provide an objective comparison of different approaches.
(a) ResNet18 with CE.
Silhouette score: 0.121
(b) ResNet18 with CDW-CE.
Silhouette score: 0.222
Figure 5: t-SNE embeddings of features obtained from ResNet18 model trained with a) CE b) CDW-CE
Figures 5(a) and 5(b) display the t-SNE plots illustrating the embeddings of our models trained on LIMUC using CE and CDW-CE, respectively. The plots reveal a notable separation between MES 0 and MES 3 classes compared to MES 1 and MES 2 classes. However, these plots do not distinctly indicate the relative embedding qualities. Therefore, we employed the Silhouette Score to quantitatively compare the feature extraction ability of the two loss functions. For $N$ samples, the average Silhouette Score is calculated as shown in Eq. 5.
$S=\frac{1}{N}\sum_{i=1}^{N}\frac{b_{i}-a_{i}}{\max(a_{i},b_{i})}$ (5)
where $a_{i}$ is the mean distance from the $i^{th}$ sample to other samples in the same cluster, and $b_{i}$ is the mean distance from the $i^{th}$ sample to samples in the nearest cluster.
The Silhouette scores for models trained with CE and CDW-CE were calculated as 0.121 and 0.222, respectively. This suggests that models trained with CDW-CE exhibit a better feature extraction ability compared to those trained with the standard CE loss function.
Explainability Analysis: Class Activation Maps
Class Activation Maps (CAM) enable the analysis of where the model focuses its attention in a given input image. In this study, we obtained CAM for two models—one trained using our CDW-CE method and the other with the CE loss function. Some sample images and the corresponding CAM visuals for both models are provided in Figure 7.
Figure 6: Expert evaluations of CAM visualizations for models trained with CE and CDW-CE. The percentage values representing the instances where experts found both visualizations equally reasonable were 37.7%, 31.1%, 57.8%, 62.8%, and 47.4%, respectively.
CAM visualizations can be employed to obtain important feedback from domain experts through a simple questionnaire, asking them to identify which model’s attention aligns more accurately with the underlying symptoms. For this task, we presented the experts with 240 images, comprising 60 images from each category. Both models correctly identified these images. The experts were only shown the original image and two CAM visualizations superimposed on the original images. The CAM images generated by the models for every new image were arbitrarily labeled as AI-1 (Artificial Intelligence-1) and AI-2. Without being aware of which model produced which CAM visualization, the clinicians were requested to choose which CAM visualization was more aligned with their decision-making. If there were no significant differences, they had the option to indicate that CAM images were equally reasonable in terms of decision-making. According to their feedback, as can be observed from Figure 6, our method enables the CNN models to better align with the expectations of the domain experts on where the model shall attend.
Figure 7: Original images and CAM visualizations of the ResNet18 model trained with CE and CDW-CE. The model trained with CDW-CE highlights broader and more relevant areas associated with the disease.
Analysis of the Hyperparameters
The power term $\alpha$ in the loss function controls the strength of the penalization, with higher values leading to stronger penalization for mispredictions that deviate further from the true class. We have integrated a margin term to the CDW-CE loss function to penalize inconsistent classifications further and make the classifier more robust regarding better classification boundaries and a more considerable distance between clusters of classes. Adding these terms aims to enhance intra-class compactness, leading to more tightly grouped in-class distributions. Similarly, the inter-class distance, or the separation between different class clusters, is intended to be improved. While this results in better classification performance and robustness, it also requires more extensive hyperparameter tuning.
The addition of the margin term improves the CDW-CE loss function in terms of classification performance as demonstrated by a ResNet18 architecture trained with and without margin term CDW-CE loss functions with different $\alpha$ values in Figure 8. Here, we have approached the margin term as a new hyperparameter and tuned it according to the QWK performance. With this tuning, we have achieved improved results for different margin values corresponding to different power terms ($\alpha$). The $\alpha=1$ is the basic CDW-CE loss function with no class distance weights, and the inclusion of the margin term improves the results in this baseline form. It is important to note that the extra hyperparameter tuning effort is required to obtain a possibly better result than default values for the margin value for different tasks and datasets.
Figure 8: CDW-CE vs. CDW-CE with margin for varying $\alpha$ values. Numbers above the blue markers show the margin value.
Adding a margin term effectively improves all three models used in this work in terms of QWK score.
Table 3: QWK scores for CDW-CE vs. CDW-CE with margin ($m$ refers to additive margin value).
Loss function ResNet18 Inception-v3 MobileNet-v3-large
CDW-CE 0.8568 0.8678 0.8588
CDW-CE w/ margin 0.8600 ($m$=0.05) 0.8719 ($m$=0.025) 0.8617 ($m$=0.0025)
Conclusions
In this work, we have proposed a new loss function, Class Distance Weighted Cross-Entropy (CDW-CE), designed specifically for ordinal classification tasks. Our experiments on the LIMUC dataset demonstrated that ordinal loss functions perform better than the widely used non-ordinal loss function CE and other ordinal loss functions. Moreover, the proposed CDW-CE method outperformed the other methods across various performance metrics and architectures. Additionally, an analysis of remission scores indicated that CDW-CE surpasses other ordinal loss functions (CO2, HO2, CORN) in remission scores.
To demonstrate the encoding capability of models trained with our approach, we generated t-SNE plots. We calculated corresponding silhouette scores for CDW-CE and CE loss functions, representing ordinal and categorical scenarios. The Silhouette Score comparison revealed that CDW-CE, with higher scores, exhibits superior intra-class grouping and better inter-class separation. Furthermore, feedback from domain experts based on CAM visualizations indicated that our model outperformed the baseline CE loss function, suggesting improved explainability.
Acknowledgements
This work has been supported by Middle East Technical University Scientific Research Projects Coordination Unit under grant number ADEP-704-2024-11486.
Data availability
Labeled Images for Ulcerative Colitis (LIMUC) [7] dataset is a publicly available dataset.
Conflicts of interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.